40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 985     Entry time: Tue Sep 23 13:25:07 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: a bit better 
I've been spending time working on the short DOF loops (PRC,MICH,SRC) in an attempt to make the
initial stage of lock acquisition (the DRMI+2ARMs, no spring) better. This seems to have been
largely successful, as last night there were several locks of the DRMI+2ARMs with pretty short
wait times.

The output matrix for the short DOFs is a bit strange, though. The MICH->PRM element is about
3 times too small, which seems to indicate something broken in hardware. The MICH->SRM element
seems normal, though, which suggests the BS is isn't broken--either the PRM has had a sudden
actuation increase or it's a problem with the sensing.
ELOG V3.1.3-